OpenStudy (anonymous):

Solve the equation: 4x2 = 36 how do i solve this?

OpenStudy (zehanz):

4x²=36. So 4 times a certain number is 36, that means that number can only be 9. But... the number is not x, it is x², so now you still need to solve: x²=9. You now have to think what number(s) is (are) 9 when squared...

OpenStudy (anonymous):

3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

it could be 3, or it could be \(-3\)

OpenStudy (zehanz):

-3

OpenStudy (zehanz):

(-3)²=(-3)*(-3)=9

OpenStudy (anonymous):

\[4x^2=36\]divide by 4 get \[x^2=9\]so \(x\) could be \(3\) since \(3^2=9\) or \(x\) could be \(-3\) because \((-3)^2=9\) too

OpenStudy (zehanz):

@satellite73: In this case I have used the "thumb"-method: put your thumb on x² and ask yourself: 4 times what is 36? This method is used to solve equations in a more informal way. Why? Because it is a simple equation and there is no need to bring out all the equipment ;) (Of course I know that you can always use your method).